# Dynamics on invariant tori emerging through forced symmetry breaking in phase oscillator networks

### v1

@@ 4.2. Perturbing $\mathrm{SDD}$
We first focus on the first-order perturbation of $\mathrm{SDD}$ for the dynamics (15) with perturbation (22) with global phase-difference coupling determined by the $2\pi$-periodic function $h$. With $i:=\sqrt{-1}$ write $h(\varphi)=\sum_{\ell\in\mathbb{Z}}h_{\ell}e^{i\ell\varphi}$ with Fourier coefficients $h_{\ell}\in\mathbb{C}$. To simplify notation, we will typically suppress the indices “$\mathrm{SDD}$” (the torus) and “$1$” (the order of the approximation in terms of the small parameter) during the computations.
@@
(33) $V_{\ell}^{12}$ $=\left(\begin{matrix}0\\ 2h_{\ell}((-1)^{\ell}-1)\\ 0\end{matrix}\right),$ $V_{\ell}^{13}$ $=\left(\begin{matrix}0\\ 0\\ 2h_{\ell}((-1)^{-\ell}-1)\end{matrix}\right).$
With (14) this yields
@@
(37a) $U_{\ell}^{00}$ $=\left(\begin{matrix}2h_{\ell}\\ (1+(-1)^{\ell})h_{\ell}\\ (1+(-1)^{\ell})h_{\ell}\end{matrix}\right),$ $U_{\ell}^{21}$ $=\left(\begin{matrix}(1+(-1)^{\ell})h_{\ell}\\ (1+(-1)^{-\ell})h_{-\ell}\\ 0\end{matrix}\right)$
(37b) $U_{\ell}^{31}$ $=\left(\begin{matrix}(1+(-1)^{\ell})h_{\ell}\\ 0\\ (1+(-1)^{-\ell})h_{-\ell}\end{matrix}\right),$ $U_{\ell}^{32}$ $=\left(\begin{matrix}0\\ (1+(-1)^{\ell})h_{\ell}\\ (1+(-1)^{-\ell})h_{-\ell}\end{matrix}\right)$
We can now choose the tangential component of the embedding $X$ such that the first order approximation $f$ is in normal form; cf. Section 2.4. Specifically, with $\langle\Omega,{\boldsymbol{\ell}}\rangle=\Omega_{1}\ell_{1}+\Omega_{2}\ell_{2}+\Omega_{3}\ell_{3}$ choose
@@
which only contains nonresonant terms. With this choice of $X$ and (14) we obtain the first order correction of the vector field on the perturbed torus
@@
We can write out the first-order approximation of the embedding and the dynamics on the perturbed torus for an explicit coupling function. First note that the nontrivial Fourier coefficients (37) and (33) vanish for odd coefficients $\ell$ of the perturbation function $h$. Thus for the biharmonic coupling function $h(\varphi)=\sin(\varphi+\alpha)+r\sin(2(\varphi+\beta))$ in Proposition 3.2 we expect that only the second harmonic appears in the reduced dynamics. Indeed, for the coupling function we have
@@ 4.3. Perturbing $\mathrm{SSD}$
$V_{\ell}^{13}$ $=\left(\begin{matrix}0\\ 0\\ 2h_{\ell}((-1)^{\ell}-1)\end{matrix}\right),$ $V_{\ell}^{23}$ $=\left(\begin{matrix}0\\ 0\\ 2h_{\ell}((-1)^{\ell}-1)\end{matrix}\right).$
This gives the normal component of the embedding,
@@
(44) $Y_{\ell}^{13}=Y_{\ell}^{23}=\left(\begin{matrix}0\\ 0\\ \left(\dfrac{2(2r_{0}-i\ell)((-1)^{\ell}-1)}{4+\ell^{2}}\right)h_{\ell}\end{matrix}\right).$
For the tangential direction, we choose
(45) $X(\phi)=\sum_{\ell\in\mathbb{Z}\smallsetminus\{0\}}\left(\frac{i}{2\ell}U_{\ell}^{31}e^{i\ell(\phi_{3}-\phi_{1})}+\frac{i}{2\ell}U_{\ell}^{32}e^{i\ell(\phi_{3}-\phi_{2})}\right)$
with
@@
(46b) $U_{\ell}^{31}$ $=\left(\begin{matrix}(1+(-1)^{\ell})h_{\ell}\\ 0\\ (1+(-1)^{-\ell})h_{-\ell}\end{matrix}\right),$ $U_{\ell}^{32}$ $=\left(\begin{matrix}0\\ (1+(-1)^{\ell})h_{\ell}\\ (1+(-1)^{-\ell})h_{-\ell}\end{matrix}\right)$
so that dynamics on the perturbed torus is in normal form. With these Fourier coefficients, we obtain
@@
so that, up to first order, the dynamics is determined by $f^{\mathrm{SSD}}_{\delta}=\Omega^{\mathrm{SSD}}+\delta f_{1}^{\mathrm{SSD}}+\mathcal{O}(\delta^{2})$ and the perturbed torus is embedded through $e^{\mathrm{SSD}}_{\delta}=E^{\mathrm{SSD}}+\delta e^{\mathrm{SSD}}_{1}+\mathcal{O}(\delta^{2})$ with $e^{\mathrm{SSD}}_{1}(\phi)=RX(\phi)+NY(\phi)$. This is the main claim of Proposition 3.1 for $\mathrm{SSD}$.
@@ 5. Numerical continuation of solutions on perturbed tori
We now explore numerically how solutions on the perturbed tori change and bifurcate as the asymmetry parameter $\delta$ is increased. The approximations computed in the previous section are the key to generate starting data for the numerical continuation: The first-order dynamics allow to compute approximate solutions for $\delta$ small on the tori $\mathrm{SDD},\mathrm{SSD}$. The embedding then allows to map these solutions into the full phase space $\mathbb{T}^{6}$ to generate starting data that can be continued using Auto (see [25] for accompanying computer code).
Note that the $S_{2}$ symmetry on $\mathbb{T}^{6}$ that acts by permuting the oscillators within one population induces a symmetry action on $\mathbb{T}^{3}$. If a population is desynchronized (‘$\mathrm{D}$’) this corresponds to a shift by $\pi$. We expect that this is reflected in the bifurcation diagrams. Moreover, this explains why a perturbation with single harmonics does not affect the local dynamics on $\mathrm{SDD}$ where only even harmonics appear in the normal form dynamics: Only even harmonics preserve a phase shift by $\pi$.
Figure 1. Dynamics of system (40) for $\alpha=\beta=\frac{\pi}{2}$, $r=0.2$ and $\delta=0.01$. Panel (a) shows solutions to the reduced system (50) on the torus $\mathbb{T}^{2}$ described by coordinates $(\phi_{2},\phi_{3})$ representing the ‘$\mathrm{D}$’ populations; these periodic orbits wind around $\mathbb{T}^{2}$. Shown are two stable periodic orbits (blue) at $\psi\in\{\frac{\pi}{2},\frac{3\pi}{2}\}$ and two unstable periodic orbits (red) at $\psi\in\{0,\pi\}$. Panel (b) shows a periodic orbit in the full system that lies on $\mathrm{SDD}$ for $\phi_{3}-\phi_{2}=0$. Note that the two ‘$\mathrm{D}$’ populations have approximately the same phase.
@@ 5.1. Dynamics on the perturbation of $\mathrm{SDD}$
As $\phi_{1}$ has constant motion, the dynamics are effectively one-dimensional; the $\psi$-nullclines correspond to periodic orbits in the full system (40), with their stability given by the equation for $\dot{\psi}$. For $\delta,r>0$, the dynamics depends on $\beta$ as follows: For $\beta^{*}=\frac{\pi}{4}+\frac{k\pi}{2},k\in\mathbb{Z},$ the torus $\mathbb{T}^{2}$ is foliated by neutrally stable periodic orbits. For $\beta\neq\beta^{*}$, the system possesses four hyperbolic periodic orbits, two of which are stable and two are unstable; this is a consequence of the second harmonics in the normal form. The two stable (resp. unstable) periodic orbits are related by a shift by $\pi$ along the corresponding population (i.e., $\phi$-coordinate) and have exactly the same Floquet multipliers (computed numerically), as expected from the $\mathbb{Z}_{2}$-symmetry translated into normal form coordinates. The stability of these periodic orbits is reversed once $\beta$ passes through $\beta^{*}$. The location of these hyperbolic periodic orbits does not change upon variations of $\beta\neq\beta^{*}$. We fix $\beta=\frac{\pi}{2}$ so that (50) has unstable periodic orbits for $\psi\in\{0,\pi\}$ and unstable periodic orbits for $\psi\in\big{\{}\frac{\pi}{2},\frac{3\pi}{2}\big{\}}$. They correspond to the periodic orbits for which $\phi_{3}-\phi_{2}\in\{0,\pi\}$ and $\phi_{3}-\phi_{2}\in\big{\{}\frac{\pi}{2},\frac{3\pi}{2}\big{\}}$, respectively; this gives an idea on how the ‘$\mathrm{D}$’ populations organize relative to each other. Figure 1 shows the organization of phase space for system (50) with $\beta=\frac{\pi}{2}$. Panel (a) shows the configuration in $(\phi_{2},\phi_{3})$-plane, relating the ‘$\mathrm{D}$’ populations. Panel (b) shows the unstable periodic orbit that corresponds to $\phi_{3}-\phi_{2}=0$ in the full system.
@@
We now want to study the dynamics on $\mathrm{SDD}$ in full space $\mathbb{T}^{6}$ for different values of $\delta$ using numerical continuation. To generate initial data, we start with an asymptotically stable periodic obit of the first-order approximation $f_{1}^{\mathrm{SSD}}$ on $\mathbb{T}^{3}$. Taking advantage of the collocation framework of Auto allows to continue these periodic orbits further. Once we obtain a good approximation of one of the stable periodic orbits corresponding to $\psi=\frac{\pi}{2}$ or $\psi=\frac{3\pi}{2}$ in system (40), we use the embedding (41) to get a periodic orbit of the perturbed system (19) on $\mathbb{T}^{6}$ that lies on a perturbation of $\mathrm{SDD}$. Note that such a periodic orbit is not stable in $\mathbb{T}^{6}$ since we carry on instabilities of the saddle torus $\mathrm{SDD}$, whose stability properties persist under small perturbations due to normal hyperbolicity. This is one of the benefits of the reduction to the perturbed tori: We directly obtain approximations for two (symmetry-related) saddle periodic orbits, which is practically impossible by standard shooting techniques. We use any of the periodic orbits obtained this way as an initial condition for studying numerically the dynamics on a perturbation of $\mathrm{SDD}$ with respect to $\delta$.
@@
Figure 2. Bifurcation diagram of periodic orbits of the full system (19) with the perturbation driven by (22) and (25) for $\delta>0$ stemming from a perturbation of $\mathrm{SDD}$. The rest of the parameters are as in (51). Panel (a) shows the four branches of hyperbolic periodic orbits $\gamma_{0},\gamma_{\frac{\pi}{2}},\gamma_{\pi}$ and $\gamma_{\frac{3\pi}{2}}$ emanating from $\psi^{*}\in\big{\{}0,\frac{\pi}{2},\pi,\frac{3\pi}{2}\big{\}}$ at $\delta=0$. The vertical green line at $\delta=0$ represents the unpertubed case, in which the torus $\mathrm{SDD}$ is foliated by neutrally stable periodic orbits. Panel (b) highlights secondary branches of periodic orbits emanating from the branches $\gamma_{0}$ and $\gamma_{\pi}$ at $\delta=\delta^{*}$ (see text) forming isolas.
@@
that specify the symmetry breaking perturbation determined by (25). Parameters (51a) are such that the torus $\mathrm{SDD}$ and $\mathrm{SSD}$ are of saddle type for $\delta=0$; see [11]. We set $\delta=0.01$ at the beginning of the computation to obtain the initial stable periodic orbit in system (40) and then embed it in $\mathbb{T}^{6}$ to start the continuation. The horizontal axis in figure 2 corresponds to the values of the perturbation parameter $\delta$, and the vertical axis corresponds to the average of the phase difference $\theta_{3,1}-\theta_{2,1}$, modulo $2\pi$, along the corresponding periodic orbit for each parameter value. We abuse notation and denote this averaged phase difference by $\psi$ as in figure 1, in order to be able to compare both figures. In fact, $\psi$ measures the relative positions between the two ‘$\mathrm{D}$’ populations in $\mathrm{SDD}$. The vertical green line in Panel (a) represents the family of neutrally stable periodic orbits that exist for $\delta=0$; they correspond to the foliation of $\mathrm{SDD}$ by periodic orbits in the fully symmetric case.
@@
Figure 3. Stable synchrony patterns bifurcating from primary branches of periodic orbits. Panels (a)–(d) show the phase evolution of oscillators in ‘$\mathrm{D}$’ populations along one period on coexisting stable periodic orbits lying on the secondary branches emanating from $\psi=0$ (bottom) and $\psi=\pi$ (top) for $\delta\approx 0.40372$; see the magnified parameter range in figure 2(b). Shading indicates the deviation from $\pi$ of each phase oscillator, where black indicates that $\theta_{\sigma,k}=\pi$ and white $\theta_{\sigma,k}=0$ or $2\pi$.
@@
We now describe the bifurcations in more detail as the asymmetry parameter $\delta$ is varied. Four branches of periodic orbits emanate for $\delta>0$ from the vertical green line in figure 2(a) at $\psi^{*}\in\big{\{}0,\frac{\pi}{2},\pi,\frac{3\pi}{2}\big{\}}$ denoted by $\gamma_{\psi^{*}}$; see the resemblance with figure 1, which corresponds to the (reduced) dynamics on phase space for a fixed value of $\delta\neq 0$ small. These four periodic orbits can be interpreted in terms of their phase configurations as invariant subsets of $\mathrm{SDD}$: In $\gamma_{0}$ the two ‘$\mathrm{D}$’ populations are phase synchronized, in $\gamma_{\frac{\pi}{2}}$ they are $\frac{\pi}{2}$ apart etc—of course this interpretation is only approximate as $\delta$ is increased. We use similar colors as for the branches of periodic orbits in figure 1, again to stress the connection between the two figures, and highlighting the stability within $\mathrm{SDD}$ for $\delta$ small (they are unstable in $\mathbb{T}^{6}$). For each $\delta>0$, the periodic orbits on $\gamma_{0}$ and $\gamma_{\pi}$ have (numerically) the same Floquet multipliers; the same occurs with the periodic orbits on $\gamma_{\frac{\pi}{2}}$ and $\gamma_{\frac{3\pi}{2}}$. This is reminiscent of the $\mathbb{Z}_{2}$ symmetry in the (unperturbed) full system that is still present in the full system for $\delta\neq 0$ small. The branches $\gamma_{0}$ and $\gamma_{\pi}$ both become stable at (numerically) the same value $\delta=\delta^{*}\approx 0.3152$, from which a secondary branch of periodic orbits emanates on each primary branch $\gamma_{0},\gamma_{\pi}$.
@@
The secondary bifurcations are shown in detail in Panel (b) of figure 2. Each of these secondary branches forms an isola of periodic orbits, and can be thought of as two secondary branches emanating from the main branch at $\delta^{*}$, that collide at the point $\mathsf{SN}_{3}$. The two secondary branches forming each isola undergo exactly the same bifurcations at (numerically) exactly the same parameter values: both (secondary) branches pass through a saddle-node ($\mathsf{SN}_{1,5}$), torus ($\mathsf{T}_{1,2}$) and another saddle-node bifurcation ($\mathsf{SN}_{2,4}$) before colliding on $\mathsf{SN}_{3}$. This suggests that the secondary branches arise from a pitchfork bifurcation whose symmetries are not present in the projection used in figure 2. This bifurcation must be transverse to the perturbed torus $\mathrm{SDD}$ (if it still exists), since all the Floquet multipliers of the corresponding periodic orbits are stable after the bifurcation. The secondary branches arising from this bifurcation cross the branch $\gamma_{\frac{3\pi}{2}}$, indicating that a torus breakdown has likely happened for smaller values of $\delta$. This is consistent with the existence of saddle-node bifurcations on the secondary branches before this crossing, which often relates to a loss of smoothness of the torus. There is a small interval of multistability between the points $\mathsf{SN}_{1}$ ($\mathsf{SN}_{5}$) and $\mathsf{T}_{1}$ ($\mathsf{T}_{2}$), where two stable periodic orbits on each secondary branch coexists with a sable periodic orbit lying on the primary branch, adding up to 6 stable periodic orbits all together. A similar structure occurs along the primary branch $\gamma_{\pi}$, but the labels are not included in figure 2(b).

### v2

@@ 4.2. Perturbing $\mathrm{SDD}$
We first focus on the first-order perturbation of $\mathrm{SDD}$ for the dynamics (16) with perturbation (19) with global phase-difference coupling determined by the $2\pi$-periodic function $h$. With $i:=\sqrt{-1}$ write $h(\varphi)=\sum_{\ell\in\mathbb{Z}}h_{\ell}e^{i\ell\varphi}$ with Fourier coefficients $h_{\ell}\in\mathbb{C}$. To simplify notation, we will typically suppress the indices $\mathrm{SDD}$” (the torus) and “11” (the order of the approximation in terms of the small parameter) during the computations.
@@
(34) $V_{\ell}^{12}$ $=\left(\begin{matrix}0\\ -4h_{\ell}\\ 0\end{matrix}\right),$ $V_{\ell}^{13}$ $=\left(\begin{matrix}0\\ 0\\ -4h_{\ell}\end{matrix}\right).$
for $\ell$ odd and and $V_{\ell}^{12}=0$ for $\ell$ even. With (14) this yields
@@
(38) $U_{\ell}^{00}$ $=\left(\begin{matrix}2h_{\ell}\\ 2h_{\ell}\\ 2h_{\ell}\end{matrix}\right),$ $U_{\ell}^{21}$ $=\left(\begin{matrix}2h_{\ell}\\ 2h_{-\ell}\\ 0\end{matrix}\right),$ $U_{\ell}^{31}$ $=\left(\begin{matrix}2h_{\ell}\\ 0\\ 2h_{-\ell}\end{matrix}\right),$ $U_{\ell}^{32}$ $=\left(\begin{matrix}0\\ 2h_{\ell}\\ 2h_{-\ell}\end{matrix}\right)$
for $\ell$ even and $U_{\ell}^{00}=\left(2h_{\ell},0,0\right)^{\mathsf{T}}$, $U_{\ell}^{21}=U_{\ell}^{31}=U_{\ell}^{32}=0$ for $\ell$ odd. We can now choose the tangential component of the embedding $X$ to contain the nonresonant terms such that the first order approximation $f$ is in normal form with only resonant terms; cf. Section 2.4. Specifically, if $\langle\Omega,{\boldsymbol{\ell}}\rangle=\Omega_{1}\ell_{1}+\Omega_{2}\ell_{2}+\Omega_{3}\ell_{3}$ then
@@
only contains nonresonant terms for which $\langle\Omega,{\boldsymbol{\ell}}\rangle\neq 0$. With this choice of $X$ and (14) we obtain the first order correction of the vector field on the perturbed torus
@@
We can write out the first-order approximation of the embedding and the dynamics on the perturbed torus for an explicit coupling function. Since the nontrivial Fourier coefficients $U_{\ell}^{21},U_{\ell}^{31}=U_{\ell}^{32}$ vanish for odd $\ell$ of the perturbation function $h$, we expect that only the second harmonic appears in the reduced dynamics for the biharmonic coupling function $h(\varphi)=\sin(\varphi+\alpha)+r\sin(2(\varphi+\beta))$ in Proposition 3.2. Indeed, for the coupling function we have
@@ 4.3. Perturbing $\mathrm{SSD}$
$V_{\ell}^{13}=V_{\ell}^{23}$ $=\left(\begin{matrix}0\\ 0\\ -4h_{\ell}\end{matrix}\right).$
for $\ell$ odd and $V_{\ell}^{13}=V_{\ell}^{23}=0$ for $\ell$ even. This gives the normal component of the embedding,
@@
(45) $Y_{\ell}^{13}=Y_{\ell}^{23}=\left(\begin{matrix}0\\ 0\\ \left(\dfrac{-4(2r_{0}-i\ell)}{4+\ell^{2}}\right)h_{\ell}\end{matrix}\right).$
for $\ell$ odd and $Y_{\ell}^{13}=Y_{\ell}^{23}=0$ for $\ell$ even.
For the tangential direction we compute the Fourier coefficients
@@
for any $\ell$ and
(46b) $U_{\ell}^{31}$ $=\left(\begin{matrix}2h_{\ell}\\ 0\\ 2h_{-\ell}\end{matrix}\right),$ $U_{\ell}^{32}$ $=\left(\begin{matrix}0\\ 2h_{\ell}\\ 2h_{-\ell}\end{matrix}\right)$
for $\ell$ even and $U_{\ell}^{31}=U_{\ell}^{32}=0$ for $\ell$ odd. Thus by choosing
(47) $X(\phi)=\sum_{\ell\in\mathbb{Z}\smallsetminus\{0\}}\left(\frac{i}{2\ell}U_{\ell}^{31}e^{i\ell(\phi_{3}-\phi_{1})}+\frac{i}{2\ell}U_{\ell}^{32}e^{i\ell(\phi_{3}-\phi_{2})}\right),$
the dynamics on the perturbed torus
@@
is in normal form. Thus, up to first order, the dynamics is determined by $f^{\mathrm{SSD}}_{\delta}=\Omega^{\mathrm{SSD}}+\delta f_{1}^{\mathrm{SSD}}+\mathcal{O}(\delta^{2})$ and the perturbed torus is embedded through $e^{\mathrm{SSD}}_{\delta}=E^{\mathrm{SSD}}+\delta e^{\mathrm{SSD}}_{1}+\mathcal{O}(\delta^{2})$ with $e^{\mathrm{SSD}}_{1}(\phi)=RX(\phi)+NY(\phi)$, which is the main claim of Proposition 3.1 for $\mathrm{SSD}$.
@@ 5. Numerical continuation of solutions on perturbed tori
We now explore numerically how solutions on the perturbed tori change and bifurcate as the asymmetry parameter $\delta$ is increased. The approximations computed in the previous section are the key to generate starting data for the numerical continuation: The first-order dynamics allow one to compute approximate solutions for $\delta$ small on the tori $\mathrm{SDD},\mathrm{SSD}$. The embedding then maps these solutions into the full phase space $\mathbb{T}^{6}$ to generate starting data that can be continued using Auto (see [25] for accompanying computer code).
Note that the $S_{2}$ symmetry on $\mathbb{T}^{6}$ that acts by permuting the oscillators within one population induces a symmetry action on $\mathbb{T}^{3}$. If a population is desynchronized ($\mathrm{D}$’) this corresponds to a shift by $\pi$. We expect that this is reflected in the bifurcation diagrams. Moreover, this explains why a perturbation with single harmonics does not affect the local dynamics on $\mathrm{SDD}$ where only even harmonics appear in the normal form dynamics: Only even harmonics preserve a phase shift by $\pi$.
Figure 1. Dynamics of system (41) for $\alpha=\beta=\frac{\pi}{2}$, $r=0.2$ and $\delta=0.01$. Panel (a) shows solutions to the reduced system (51) on the torus $\mathbb{T}^{2}$ described by coordinates $(\phi_{2},\phi_{3})$ representing the $\mathrm{D}$’ populations; these periodic orbits wind around $\mathbb{T}^{2}$. Shown are two stable periodic orbits (blue) at $\psi\in\{\frac{\pi}{2},\frac{3\pi}{2}\}$ and two unstable periodic orbits (red) at $\psi\in\{0,\pi\}$. Panel (b) shows a periodic orbit in the full system that lies on $\mathrm{SDD}$ for $\phi_{3}-\phi_{2}=0$. Note that the two $\mathrm{D}$’ populations have approximately the same phase.
@@ 5.1. Dynamics on the perturbation of $\mathrm{SDD}$
As $\phi_{1}$ has constant motion, the dynamics are effectively one-dimensional; the $\psi$-nullclines correspond to periodic orbits in the full system (41), with their stability given by the equation for $\dot{\psi}$. For $\delta,r>0$, the dynamics depends on $\beta$ as follows: For $\beta^{*}=\frac{\pi}{4}+\frac{k\pi}{2},k\in\mathbb{Z},$ the torus $\mathbb{T}^{2}$ is foliated by neutrally stable periodic orbits. For $\beta\neq\beta^{*}$, the system possesses four hyperbolic periodic orbits, two of which are stable and two are unstable; this is a consequence of the second harmonics in the normal form. The two stable (resp. unstable) periodic orbits are related by a shift by $\pi$ along the corresponding population (i.e., $\phi$-coordinate) and have exactly the same Floquet multipliers (computed numerically), as expected from the $\mathbb{Z}_{2}$-symmetry translated into normal form coordinates. The stability of these periodic orbits is reversed once $\beta$ passes through $\beta^{*}$. The location of these hyperbolic periodic orbits does not change upon variations of $\beta\neq\beta^{*}$. We fix $\beta=\frac{\pi}{2}$ so that (51) has unstable periodic orbits for $\psi\in\{0,\pi\}$ and stable periodic orbits for $\psi\in\big\{\frac{\pi}{2},\frac{3\pi}{2}\big\}$. They correspond to the periodic orbits for which $\phi_{3}-\phi_{2}\in\{0,\pi\}$ and $\phi_{3}-\phi_{2}\in\big\{\frac{\pi}{2},\frac{3\pi}{2}\big\}$, respectively; this gives an idea on how the $\mathrm{D}$’ populations organize relative to each other. Figure 1 shows the organization of phase space for system (51) with $\beta=\frac{\pi}{2}$. Panel (a) shows the configuration in $(\phi_{2},\phi_{3})$-plane, relating the $\mathrm{D}$’ populations. Panel (b) shows the unstable periodic orbit that corresponds to $\phi_{3}-\phi_{2}=0$ in the full system.
@@
We now want to study the dynamics on $\mathrm{SDD}$ in full space $\mathbb{T}^{6}$ for different values of $\delta$ using numerical continuation. To generate initial data, we start with an asymptotically stable periodic obit of the first-order approximation $f_{1}^{\mathrm{SSD}}$ on $\mathbb{T}^{3}$. These periodic orbits can then be continued by taking advantage of the collocation framework of Auto. Once we obtain a good approximation of one of the stable periodic orbits corresponding to $\psi=\frac{\pi}{2}$ or $\psi=\frac{3\pi}{2}$ in system (41), we use the embedding (42) to get a periodic orbit of the perturbed system (16) on $\mathbb{T}^{6}$ that lies on a perturbation of $\mathrm{SDD}$. Note that such a periodic orbit is not stable in $\mathbb{T}^{6}$ since we carry on instabilities of the saddle torus $\mathrm{SDD}$, whose stability properties persist under small perturbations due to normal hyperbolicity. This is one of the benefits of the reduction to the perturbed tori: We directly obtain approximations for two (symmetry-related) saddle periodic orbits, which is practically impossible by standard shooting techniques. We use any of the periodic orbits obtained this way as an initial condition for studying numerically the dynamics on a perturbation of $\mathrm{SDD}$ with respect to $\delta$.
@@
Figure 2. Bifurcations can stabilize periodic orbits of the full system (16) stemming from $\mathrm{SDD}$ for a perturbation determined by (19) and (26) that is sufficiently large $\delta>0$. The rest of the parameters are as in (52). Panel (a) shows the four branches of hyperbolic periodic orbits $\gamma_{0},\gamma_{\frac{\pi}{2}},\gamma_{\pi}$ and $\gamma_{\frac{3\pi}{2}}$ emanating from $\psi^{*}\in\big\{0,\frac{\pi}{2},\pi,\frac{3\pi}{2}\big\}$ at $\delta=0$, bifurcations that change their stability in the full system, and secondary solution branches. The vertical green line at $\delta=0$ represents the unperturbed case, in which the torus $\mathrm{SDD}$ is foliated by neutrally stable periodic orbits. Panel (b) highlights bifurcations at $\delta=\delta^{*}$ that stabilize the branches $\gamma_{0}$ and $\gamma_{\pi}$ in the full system and corresponding secondary branches of periodic orbits forming isolas.
@@
that specify the symmetry breaking perturbation determined by (26). Parameters (52a) are such that the torus $\mathrm{SDD}$ and $\mathrm{SSD}$ are of saddle type for $\delta=0$; see [11]. We set $\delta=0.01$ at the beginning of the computation to obtain the initial stable periodic orbit in system (41) and then embed it in $\mathbb{T}^{6}$ to start the continuation. The horizontal axis in Figure 2 corresponds to the values of the perturbation parameter $\delta$, and the vertical axis corresponds to the average of the phase difference $\theta_{3,1}-\theta_{2,1}$, modulo $2\pi$, along the corresponding periodic orbit for each parameter value. We abuse notation and denote this averaged phase difference by $\psi$ as in Figure 1, in order to be able to compare both figures. In fact, $\psi$ measures the relative positions between the two $\mathrm{D}$’ populations in $\mathrm{SDD}$. The vertical green line in Panel (a) represents the family of neutrally stable periodic orbits that exist for $\delta=0$; they correspond to the foliation of $\mathrm{SDD}$ by periodic orbits in the fully symmetric case.
@@
Figure 3. Stable synchrony patterns bifurcating from primary branches of periodic orbits. Panels (a)–(d) show the phase evolution of oscillators in $\mathrm{D}$’ populations along one period on coexisting stable periodic orbits lying on the secondary branches emanating from $\psi=0$ (bottom) and $\psi=\pi$ (top) for $\delta\approx 0.40372$; see the magnified parameter range in Figure 2(b). Shading indicates the deviation from $\pi$ of each phase oscillator, where black indicates that $\theta_{\sigma,k}=\pi$ and white $\theta_{\sigma,k}=0$ or $2\pi$.
@@
As the asymmetry parameter $\delta$ is varied, four branches of periodic orbits emanate for $\delta>0$ from the vertical green line in Figure 2(a) at $\psi^{*}\in\big\{0,\frac{\pi}{2},\pi,\frac{3\pi}{2}\big\}$ denoted by $\gamma_{\psi^{*}}$; see the resemblance with Figure 1, which corresponds to the (reduced) dynamics on phase space for a fixed value of $\delta\neq 0$ small. These four periodic orbits can be interpreted in terms of their phase configurations as invariant subsets of $\mathrm{SDD}$: In $\gamma_{0}$ the two $\mathrm{D}$’ populations are phase synchronized, in $\gamma_{\frac{\pi}{2}}$ they are $\frac{\pi}{2}$ apart etc—of course this interpretation is only approximate as $\delta$ is increased. We use similar colors as for the branches of periodic orbits in Figure 1, again to stress the connection between the two figures, and highlighting the stability within $\mathrm{SDD}$ for $\delta$ small (they are unstable in $\mathbb{T}^{6}$). For each $\delta>0$, the periodic orbits on $\gamma_{0}$ and $\gamma_{\pi}$ have (numerically) the same Floquet multipliers; the same occurs with the periodic orbits on $\gamma_{\frac{\pi}{2}}$ and $\gamma_{\frac{3\pi}{2}}$. This is reminiscent of the $\mathbb{Z}_{2}$ symmetry in the (unperturbed) full system that is still present in the full system for $\delta\neq 0$ small.
@@
Note that the branches $\gamma_{0}$ and $\gamma_{\pi}$ both become stable at (numerically) the same value $\delta=\delta^{*}\approx 0.3152$, from which a secondary branch of periodic orbits emanates on each primary branch $\gamma_{0},\gamma_{\pi}$. The secondary bifurcations are shown in detail in Panel (b) of Figure 2. Each of these secondary branches forms an isola of periodic orbits, and can be thought of as two secondary branches emanating from the main branch at $\delta^{*}$, that collide at the point $\mathsf{SN}_{3}$. The two secondary branches forming each isola undergo exactly the same bifurcations at (numerically) exactly the same parameter values: both (secondary) branches pass through a saddle-node ($\mathsf{SN}_{1,5}$), torus ($\mathsf{T}_{1,2}$) and another saddle-node bifurcation ($\mathsf{SN}_{2,4}$) before colliding on $\mathsf{SN}_{3}$. This suggests that the secondary branches arise from a pitchfork-like bifurcation whose symmetries are not present in the projection used in Figure 2. This bifurcation must be transverse to the perturbed torus $\mathrm{SDD}$ (if it still exists), since all the Floquet multipliers of the corresponding periodic orbits are stable after the bifurcation. This suggests that the pitchfork-like scenario is combined with a different phenomenom. In fact, the secondary branches arising from this bifurcation cross the branch $\gamma_{\frac{3\pi}{2}}$, indicating that a torus breakdown has likely happened for smaller values of $\delta$. This is consistent with the existence of saddle-node bifurcations on the secondary branches before this crossing, which often relates to a loss of smoothness of the torus. A deeper exploration of the bifurcations involving the torus breakdown is beyond of the scope of this paper, as we focus on what emanates from this primary bifurcation scenario. There is a small interval of multistability between the points $\mathsf{SN}_{1}$ ($\mathsf{SN}_{5}$) and $\mathsf{T}_{1}$ ($\mathsf{T}_{2}$), where two stable periodic orbits on each secondary branch coexists with a stable periodic orbit lying on the primary branch, adding up to 6 stable periodic orbits all together. A similar structure occurs along the primary branch $\gamma_{\pi}$, but the labels are not included in Figure 2(b).